in triangle ABC a=5,b=4,A=90+B ,then FIND VALUE OF angle C
 
 

Grade:9

1 Answers

Saurabh Koranglekar
4 years ago
Dear Student

Using sin A/ a = sin B /b

- cos B / 5 = sin B / 4

Cos B = -5 sqr( k )

Sin B = 4 sqr( k)

Sqr( k) = 1/ ( 16+25)

= 1/ 41

Using k we can identify cos B or sin B then sin A and finally sin C

Regards